OpenStudy (anonymous):

If y varies directly with x and y = -7 when x = -1, find x when y = -84. Answer 84 7 -12

OpenStudy (anonymous):

try dividing 84 by 7 to find the ratio

OpenStudy (anonymous):

-12?

OpenStudy (anonymous):

-84 divided by -7 equals +12 +12 is your ratio you need to put the x value you know into the ratio -1 times 12 equals -12 so thats the answer but you need to remember to use the ratio to find it otherwise you can end up with the wrong sign or if x had been 2 you would have a completly different answer.

OpenStudy (anonymous):

Write the slope-intercept form of the line which passes through (4, 2) with a slope of 1/2. Answer y = 1/2 x + 2 y = 1/2 x y = 1/2 x + 4